Durbin Corporation reported net sales of $259,300, cost of goods sold of $135,900, operating expenses of $48,290, net income of $42,120, beginning total assets of $514,600, and ending total assets of $572,000.

Calculate profit margin and gross profit rate. (Round answers to 0 decimal places, e.g. 10%.)
Answer #1
Concepts and reason

Accounting: Accounting is a process of recording the transactions, classifying them in a specific manner, and then it is the process of summarizing, analyzing, and interpreting the results. It is a process of preserving the accounts.

Financial accounting is a process of preparing reports to provide all financial informations to both the internal and external users of an organization. The financial statements that are prepared under the financial accounting are examined by independent certified public accountants at the year-end, who would express their opinions on the fairness of the reports shown by the company.

Income statement: This is the financial statement of a company which reports all the revenues that are earned and expenses that are to be expended by the company on the immediate accounting year. Income statement is also known profit and loss statement

Ratio analysis: Ratio analysis is an analysis used for analyzing various performances and progresses of the company. There are many ratios such as profitability, solvency, liquidity, debt–service coverage, and proprietary for analyzing.

Fundamentals

Gross Profit: Gross profit is the excess amount of sales price over estimated costs. It is the profit after deducting the manufacturing costs of product. Thus, it is the net of sales and cost of goods sold.

Net income: This is the income of a company which is the result of all the revenues earned and expenses incurred by the company over a specific period of time. It is the income after adjusting the profit and losses. It is also known as earnings after interest and taxes.

Cost of goods sold: The cost that is incurred by a business to sell products in a particular period is called cost of goods sold. It is also known as cost of sales. It is considered as the expense of current period.

Sales: Sales is an activity of selling the goods in the market which is sold by a seller and purchased by a buyer. It is the main source of revenue for the company. It is necessary to have consideration for sales.

Operating expenses: The expenses which are incurred in day-to-day business activities but not directly allied with the production of goods and service are called operating expenses. This expense is considered as a main element in the calculation for operating income.

Assets: Asset is the resource of a company to generate income. Asset is generally classified into fixed assets, current assets, tangible assets, and intangible assets. Fixed assets are the assets that are used to generate income over a long period. Current assets are the assets that are realized within the current financial year. Tangible assets are the assets that can be felt and touched. Intangible assets are the assets that cannot be felt or touched.

Determine the gross profit:

Grossprofit=SalesCostofgoodssold=$259,300$135,900=$123,400\begin{array}{c}\\{\rm{Gross profit = Sales}} - {\rm{Cost of goods sold}}\\\\{\rm{ = \$ 259,300}} - \$ 135,900\\\\ = \$ 123,400\\\end{array}

Therefore, the gross profit is $123,400.

Calculate the profit margin and gross profit rate.

Profitmargin=NetincomeSales=$42,120$259,300=16%Grossprofitrate=GrossprofitSales=$123,400$259,300=48%\begin{array}{c}\\{\rm{Profit margin = }}\frac{{{\rm{Net income}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 42,120}}{{\$ 259,300}}\\\\ = 16\% \\\\{\rm{Gross profit rate}} = \frac{{{\rm{Gross profit}}}}{{{\rm{Sales}}}}\\\\ = \frac{{\$ 123,400}}{{\$ 259,300}}\\\\ = 48\% \\\end{array}

Therefore, the profit margin is 16 percent and the gross profit rate is 48 percent.

Ans:

The profit margin is 16 percent and the gross profit rate is 48 percent.